# Equivariant Hilbert and Ehrhart series under translative group actions

### v2

@@ Abstract.
We study the equivariant Hilbert series of the complex Stanley–Reisner ring of a simplicial complex $\Sigma$ under the action of a finite group via simplicial automorphisms. We prove that, when $\Sigma$ is Cohen–Macaulay and the group action is translative, i.e., color-preserving for some proper coloring of $\Sigma$, the equivariant Hilbert series of $\Sigma$ admits a rational expression whose numerator is a positive integer combination of irreducible characters.
We relate such a result to equivariant Ehrhart theory via an equivariant analogue of the Betke–McMullen theorem developed by Jochemko, Katthän and Reiner: namely, if a lattice polytope admits a lattice unimodular triangulation that is invariant under the action of a finite group, then the equivariant Ehrhart series of the polytope equals the equivariant Hilbert series of such a triangulation.
@@ 1. Introduction
A useful result first stated by Betke and McMullen [BM85] establishes a link between the Ehrhart series of a lattice polytope and the Hilbert series of any of its unimodular lattice triangulations.
Theorem (Betke–McMullen).
@@
Now assume that a given lattice polytope, respectively simplicial complex, carries an action of a finite group $G$. We would like to modify the generating functions in (1) and (2) so that they can take care of the symmetries at play. This can be done by turning the coefficients – formerly nonnegative integers – into characters: more precisely, effective elements of the ring $\operatorname{R}_{G}$ of (complex) virtual characters of the group, see Section 3.1. This gives rise to equivariant versions of Ehrhart and Hilbert series: see Definition 5.1 and Section 3.2.
The foundations of equivariant Ehrhart theory have been laid by Stapledon in [Sta11]. Since then, several particular instances have been investigated ([ASVM20], [ASVM21], [CHK23]), with a special focus on the permutahedron. Among the publications spurred by Stapledon’s work, let us mention the recent work of Elia, Kim and Supina [EKS23], where the theory was further developed with an eye towards methods for computing equivariant Ehrhart series by finding subdivisions of the given polytope that are in various ways compatible with the given group actions. We mention here that, as the present paper was in preparation, a preprint by Stapledon [Sta23] containing new general results and techniques appeared on arXiv. After some friendly exchanges, the updated versions of both our paper and Stapledon’s make use of each other’s results.
The study of symmetries of simplicial complexes – especially order complexes of posets – has a long history, see for instance [Sta82] and [GS84]. Our approach to this topic is influenced by work of Stembridge [Ste94] studying the setting where a group acts on the simplicial complex arising from a complete fan, in view of applications to crystallographic root systems. In the recent article [AR23], Adams and Reiner have studied the equivariant Hilbert series of an abstract simplicial complex with particular focus on the case where the group action preserves a proper coloring (see Definition 2.1) of the complex itself. Their work notably contains a “colorful” equivariant version of the celebrated result by Hochster expressing graded Betti numbers of the Stanley–Reisner ring of a simplicial complex in terms of homologies of certain subcomplexes.
Our contributions mostly revolve around a special kind of group action called translative (see Definition 2.2 for the most general setting). A finite group $G$ acts translatively on a simplicial complex $\Sigma$ precisely when there exists a proper coloring of $\Sigma$ where all vertices in any given $G$-orbit share the same color. In other words, the action of $G$ on $\Sigma$ is translative if and only if it is color-preserving for some proper coloring of $\Sigma$. Proposition 4.6 highlights a strong connection between translativity and commutative algebra: the action of $G$ on $\Sigma$ is translative if and only if the Stanley–Reisner ring of $\Sigma$ (over any infinite field $\mathbb{K}$) admits a linear system of parameters entirely contained in the invariant subring $\mathbb{K}[\Sigma]^{G}$. As a consequence, when $\Sigma$ is a Cohen–Macaulay complex and $G$ acts translatively on it, the equivariant Hilbert series of $\Sigma$ has a rational expression whose numerator is a polynomial with effective coefficients, i.e., each coefficient is a positive integer combination of irreducible $G$-characters (Theorem 4.8). When $\Sigma$ is Cohen–Macaulay and admits a balanced coloring that is preserved by the action of $G$, then an even finer statement is available (Lemma 4.15), as already noted by Stanley [Sta82] for order complexes of Cohen–Macaulay posets (compare [AR23, p. 157]).
An equivariant version of the Betke–McMullen theorem cited earlier enables us to relate the previous results to Ehrhart theory. We summarize our findings in the following theorem, pointing the reader to Section 4.1 and Section 4.2 for the definition of $\mathfrak{h}_{i}$ and $\mathfrak{h}_{S}$:
Theorem (see Theorem 5.2 below).
@@
An earlier version of this manuscript proved that (3) holds under some additional hypothesis on the action $\rho$, see Remark 5.5. We subsequently learned that the more general version stated above was proved in unpublished notes by Katharina Jochemko, Lukas Katthän and Victor Reiner. We thank them for sharing their unpublished work and allowing us to include their proof. Moreover, in the final stages of preparation of this manuscript, we became aware that Alan Stapledon independently obtained an even more general form of (3), see [Sta23, Proposition 4.40, Remark 4.41]. See Remark 5.4.
Even in its strongest form, Theorem 5.2 applies to some natural and important examples of polytopes and affords explicit computations. A notable example is given by order polytopes of posets and, more generally, alcoved polytopes.
The order polytope $\mathscr{O}(X)$ associated with a partially ordered set $X$ has been defined by Stanley [Sta86] and has received wide attention ever since. We show that every poset automorphism of $X$ induces a translative action on the canonical regular unimodular triangulation of $\mathscr{O}(X)$, and such an action preserves a balanced coloring of the triangulation (Lemma 6.4). We give a formula for the equivariant Ehrhart series of $\mathscr{O}(X)$ and we illustrate our result by some explicit computations (see Example 6.9 and Proposition 6.10).
Alcoved polytopes have been introduced by Lam and Postnikov [LP07] as $\mathbb{Z}^{d}$-lattice polytopes that are triangulated by a certain standard simplicial subdivision of $\mathbb{R}^{d}$ (see Equation 24). The same authors in [LP18] later extended the theory to polytopes that are triangulated by the “alcoves” of any affine Coxeter system (see Definition 7.2 for a precise definition), with the previous case corresponding to the Coxeter case $A_{d}$ via an affine isomorphism that we describe in Remark 7.7. In Section 7 we show that, if $P$ is an alcoved polytope that is invariant by the action of some subgroup of the associated affine Coxeter group, then the induced action on the alcoved triangulation of $P$ satisfies the most restrictive hypotheses of Theorem 5.2 and thus, for instance, has effective equivariant $h$- and flag $h$-vector (see Theorem 7.6). Note that, even though order polytopes are a special case of alcoved polytopes (see Remark 7.9), we decided to treat the former separately for the benefit of those readers not familiar with the Coxeter viewpoint.
@@
Our paper is organized as follows. Section 2 contains our setup on lattice polytopes and triangulations and defines translative and proper group actions. Section 3 introduces the concept of equivariant Hilbert series, while Section 4 specializes it to the context of Stanley–Reisner rings, with a focus on the case of translative actions on Cohen–Macaulay complexes. In Section 5 we recall some basics of equivariant Ehrhart theory and relate the equivariant Ehrhart series of a lattice polytope and the equivariant Hilbert series of a unimodular triangulation under suitable hypotheses. Section 6 is devoted to the study of the equivariant Ehrhart series of order polytopes under an action of a finite group via poset automorphisms. Finally, in Section 7 we apply our results to alcoved polytopes and Lipschitz polytopes of posets.
Acknowledgements. We thank Alan Stapledon for friendly discussions on his latest work [Sta23] and catching some mistakes in an earlier draft, Victor Reiner for sharing his unpublished work with Katharina Jochemko and Lukas Katthän, and Matt Beck for his valuable comments. We also thank Mariel Supina for introducing us to the topic with her talk at the workshop Combinatorial and Algebraic Aspects on Lattice Polytopes at Kwansei Gakuin University, and we are grateful to the organizers of this workshop for their support. AD is a member of INdAM–GNSAGA and has been partially supported by the PRIN 2020355B8Y grant “Squarefree Gröbner degenerations, special varieties and related topics” and by the INdAM–GNSAGA project “New theoretical perspectives via Gröbner bases”, CUP E53C22001930001.
@@ 2.3. Translative and proper actions
Translative actions on posets arose from the study of matroid structures related to toric arrangements [DR18], and were further investigated in [DD21] (in the context of invariant rings of Stanley–Reisner rings) and in [BD22] (in connection to supersolvability of geometric posets). If $X$ is the poset of faces of a simplicial fan, proper actions are Stembridge’s proper actions [Ste94]. If $X$ is any poset regarded as a small category without loops, any group action in the sense of Bridson and Haefliger [BH99, Definition 1.11] is proper.
@@ 3.1. Representations of finite groups
Let $G$ be a finite group and $\mathbb{K}$ be a field. A (finite-dimensional) representation of $G$ over $\mathbb{K}$ is any homomorphism $\rho\colon G\to\mathrm{GL}(V)$, where $V$ is a finite-dimensional $\mathbb{K}$-vector space. A morphism between $\rho_{1}\colon G\to\mathrm{GL}(V_{1})$ and $\rho_{2}\colon G\to\mathrm{GL}(V_{2})$ is any $\mathbb{K}$-linear map $f\colon V_{1}\to V_{2}$ such that $f\circ\rho_{1}(g)=\rho_{2}(g)\circ f$ for every $g\in G$. With every representation $\rho$ is associated a character $\chi_{\rho}:G\mapsto\mathbb{K}$ defined by $\chi_{\rho}(g)=\operatorname{trace}(\rho(g))$. Every character is an element of the ring $\operatorname{C}_{\mathbb{K}}(G)$ of class functions (i.e., $\mathbb{K}$-valued functions on $G$ that are constant along conjugation classes of $G$). The ring of virtual characters $\operatorname{R}_{G}(\mathbb{K})$ is the subring of $\operatorname{C}_{\mathbb{K}}(G)$ generated by the characters.
The representation ring $R(G,\mathbb{K})$ (or “Grothendieck ring of representations”) of $G$ over $\mathbb{K}$ is the ring obtained by the following procedure (see, e.g., [AR23, Definition 2.3]): take the free abelian group generated by isomorphism classes of finite-dimensional representations of $G$ over $\mathbb{K}$ and quotient it by the $\mathbb{Z}$-span of the relations $[\rho_{2}]-([\rho_{1}]+[\rho_{3}])$, where $0\to\rho_{1}\to\rho_{2}\to\rho_{3}\to 0$ is a short exact sequence of representations (in particular, this implies that $[\rho]+[\rho^{\prime}]=[\rho\oplus\rho^{\prime}]$ in $R(G,\mathbb{K})$). Then extend to the whole ring the multiplication given by the tensor product of representations, i.e., $[\rho]\cdot[\rho^{\prime}]=[\rho\otimes\rho^{\prime}]$. The function $\rho\mapsto\chi_{\rho}$ maps $R(G,\mathbb{K})$ to $\operatorname{R}_{G}(\mathbb{K})$. When $\mathbb{K}$ has characteristic zero this mapping is injective and defines an isomorphism onto its image (see, e.g., [CR06, Corollary 30.14]).
@@ 3.2. The equivariant Hilbert series of a Cohen–Macaulay standard graded algebra
where $[R_{i}]$ is the isomorphism class of the $G$-representation on $R_{i}$ induced by $\rho$, as an equivariant analogue of the Hilbert series of $R$. Since for the scope of this paper we are only interested in complex representations of finite groups, unless otherwise stated we will assume that $G$ is finite, $\mathbb{K}=\mathbb{C}$, and reserve the name equivariant Hilbert series of $R$ for the following formal power series in $\operatorname{R}_{G}[[t]]$:
@@
where $\chi_{R_{i}}$ is the (isomorphism class of the) character of the complex $G$-representation of $R_{i}$ induced by $\rho$.
In the non-equivariant case (over an arbitrary field $\mathbb{K})$, it is customary to express the Hilbert series rationally as $\frac{h(t)}{(1-t)^{d}}$, where $h(t)$ is a polynomial with integer coefficients and $d$ is the Krull dimension of $R$. Such a rational expression is unique and $h(t)$ is known as the $h$-polynomial of $R$. However, when dealing with equivariant Hilbert series, things become more complicated. In the special case when $R$ is the Stanley–Reisner ring of a simplicial complex endowed with a proper action (in the sense of Definition 2.2), we will see in Section 4.1 how to use a suitable multigraded structure of $R$ to obtain a rational expression of the equivariant Hilbert series with a power of $(1-t)$ as its denominator. Even in that case, though, other rational expressions are possible, and it is not a priori obvious which choice of a denominator is the most informative one.
Recall from commutative algebra that a homogeneous system of parameters for the $\mathbb{N}$-graded $d$-dimensional $\mathbb{K}$-algebra $R$ is a collection $\theta_{1},\ldots,\theta_{d}$ of elements of $\mathfrak{m}=(x\in R\mid\deg(x)>0)$ such that $\sqrt{(\theta_{1},\ldots,\theta_{d})}=\mathfrak{m}$. If $\theta_{1},\ldots,\theta_{d}$ all have the same degree $r$, we say that the homogeneous system of parameters has degree $r$; in the special case when $r=1$, we say that the linear forms $\theta_{1},\ldots,\theta_{d}$ form a linear system of parameters for $R$. When the ring $R$ is Cohen–Macaulay, then every homogeneous system of parameters is a regular sequence (and this is in fact an alternative characterization of Cohen–Macaulayness for $R$: see, e.g., [Sta96, Theorem I.5.9]).
@@
Now take $\ell^{\prime}=x+y$ and $J^{\prime}=(\ell^{\prime})$. Now the denominator is easier to compute; indeed, since $\ell^{\prime}$ is fixed by the $\mathbb{Z}_{2}$-action, we just get that $\det(I-J^{\prime}_{1}t)=1-t$. Since the nontrivial element of $\mathbb{Z}_{2}$ sends $\overline{x}\in R/J^{\prime}$ to $-\overline{x}$, the equivariant Hilbert series of $R/J^{\prime}$ is $1+\chi_{\operatorname{sgn}}t$, and overall we have that
@@
This example, despite its simplicity, is quite instructive: indeed, {}=$\{\ell^{\prime}\}=\{x+y\}$ is a linear system of parameters for $R$ that is contained in the invariant subring $R^{\mathbb{Z}_{2}}$. We will see in Proposition 4.6 a characterization of when such linear systems of parameters can be constructed for Stanley–Reisner rings. On the other hand, choosing the linear system of parameters {}=$\{\ell\}=\{x-y\}$ yields a palindromic numerator in the rational expression of the equivariant Hilbert series of $R$, mirroring what happens for the non-equivariant Hilbert series of the Gorenstein ring $R$.

### v3

@@ Abstract.
We study representations of finite groups on Stanley–Reisner rings of simplicial complexes and on lattice points in lattice polytopes. The framework of translative group actions allows us to use the theory of proper colorings of simplicial complexes without requiring an explicit coloring to be given.
We prove that the equivariant Hilbert series of a Cohen–Macaulay simplicial complex under a translative group action admits a rational expression whose numerator is a positive integer combination of irreducible characters. This implies an analogous rational expression for the equivariant Ehrhart series of a lattice polytope with a unimodular triangulation that is invariant under a translative group action.
@@ 1. Introduction
The Ehrhart series of a lattice polytope and the Hilbert series of any of its unimodular lattice triangulations (if one exists) are linked as follows.
Theorem (Betke–McMullen [BM85]).
@@
If a group $G$ acts on the given lattice polytope, respectively simplicial complex, equivariant analogues of the generating functions in (1) and (2) are defined by replacing their coefficients by effective elements of the ring $\operatorname{R}_{G}$ of (complex) virtual characters of the group, see Section 3.1, Definition 5.1 and Section 3.2.
This point of view led Stapledon to lay the foundations of equivariant Ehrhart theory in [Sta11]. Since then, this topic has been investigated by several authors, see [ASVM20], [ASVM21], [CHK23]. In particular, Elia, Kim and Supina [EKS24] developed methods for computing equivariant Ehrhart series via special subdivisions of the given polytope. The object of our paper is an equivariant version of the above-mentioned Betke-McMullen theorem (see the Disclaimer below) and its implications for the structure of the equivariant analogues of Ehrhart and Hilbert series. Our main result in this sense is B below, which we obtain as a consequence of a more general statement about group actions on simplicial complexes.
The study of group actions on posets and simplicial complexes is a classical topic, see for instance work of Stanley [Sta82], Garsia and Stanton [GS84]. In particular, Stembridge [Ste94] introduced the notion of proper group actions on simplicial complexes. More recently, Adams and Reiner [AR23] have studied the equivariant Hilbert series of an abstract simplicial complex, focusing in particular on the case where the group action preserves a proper coloring (see Definition 2.1) of the complex itself.
An action of a group $G$ on a partially ordered set is called translative if it satisfies a simple local condition (see Definition 2.2). The word “translative” was introduced in [DR18] in the context of hyperplane arrangements that are preserved by a group of translations, as an abstraction of the induced action on the arrangement’s poset of intersections. Such group actions have appeared in various contexts in the literature, see Remark 2.3. In the case of posets of faces of a simplicial complex, we observe that translative actions coincide with actions that preserve some proper coloring of the complex (Lemma 2.4). Moreover, translativity is a strictly stronger condition than properness in the sense of Stembridge (Lemma 2.7 and Remark 2.8). Our first main result is the following characterization of translativity in terms of linear systems of parameters.
Theorem A (see Theorem 4.6 below).
Let $\mathbb{K}$ be an infinite field. The action of a group $G$ on a finite nonempty simplicial complex $\Sigma$ is translative if and only if the ring $\mathbb{K}[\Sigma]$ has a linear system of parameters $\{\theta_{i}\}_{i}$ where each $\theta_{i}$ lies in the invariant subring $\mathbb{K}[\Sigma]^{G}$.
As a consequence, when $\Sigma$ is a Cohen–Macaulay complex and $G$ acts translatively on it, the equivariant Hilbert series of $\Sigma$ has a rational expression whose numerator is a polynomial with effective coefficients (Theorem 4.8). An equivariant version of the aforementioned Betke–McMullen theorem enables us to translate these considerations into the context of Ehrhart theory, obtaining the following result (see Section 4.1 and Section 4.2 for the definition of $\mathfrak{h}_{i}$ and $\mathfrak{h}_{S}$).
Theorem B (see Theorem 5.2 below).
@@
Even in its most restrictive form, Theorem 5.2 applies to some natural and important examples of polytopes and affords explicit computations. A notable example is given by order polytopes of posets and, more generally, alcoved polytopes.
The order polytope $\mathscr{O}(X)$ associated with a partially ordered set $X$ has been defined by Stanley [Sta86] and has received wide attention ever since. We show that every poset automorphism of $X$ induces a translative action on the canonical regular unimodular triangulation of $\mathscr{O}(X)$ (see Section 6), and such an action preserves a balanced coloring of the triangulation (Lemma 6.4). We give a formula for the equivariant Ehrhart series of $\mathscr{O}(X)$ and we illustrate our result by some explicit computations (see Example 6.9 and Proposition 6.10).
Theorem C (see Theorem 6.7 below for a more detailed statement).
Let $G$ be a group of automorphisms of a finite poset $X$. Then the equivariant Ehrhart series of $\mathscr{O}(X)$ coincides with the equivariant Hilbert series of its canonical triangulation. Moreover, all entries of the equivariant flag $h$-vector are effective.
Alcoved polytopes have been introduced by Lam and Postnikov [LP07] as $\mathbb{Z}^{d}$-lattice polytopes that are triangulated by a certain standard simplicial subdivision of $\mathbb{R}^{d}$ (see Equation 24). The same authors in [LP18] later extended the theory to polytopes that are triangulated by the “alcoves” of any affine Coxeter system (see Definition 7.2 for a precise definition), with the previous case corresponding to the Coxeter case $A_{d}$ via an affine isomorphism that we describe in Remark 7.7. In Section 7 we show that, if $P$ is an alcoved polytope that is invariant by the action of some subgroup of the associated affine Coxeter group, then the induced action on the alcoved triangulation of $P$ satisfies the most restrictive hypotheses of Theorem 5.2 and thus, for instance, has effective equivariant $h$- and flag $h$-vectors.
Theorem D (see Theorem 7.6 below for a more detailed statement).
Let $P$ be an alcoved polytope of type $W$ and let $G$ be a subgroup of $\widetilde{W}$ preserving $P$. Then the equivariant Ehrhart series of $P$ equals the equivariant Hilbert series of the alcoved triangulation of $P$ (with respect to the induced $G$-action). Morever, all entries of the equivariant flag $h$-vector are effective.
Even though order polytopes are a special case of alcoved polytopes (see Remark 7.9), we decided to treat the former separately for the benefit of those readers not familiar with the Coxeter viewpoint.
@@
Our paper is organized as follows. In Section 2 we lay out our setup on lattice polytopes and triangulations, we define translative and proper group actions and relate them to other notions of group actions in the literature. Section 3 introduces the concept of equivariant Hilbert series, while Section 4 specializes it to the context of Stanley–Reisner rings, proving A and its consequences for the case of Cohen–Macaulay complexes. In Section 5 we recall some basics of equivariant Ehrhart theory and relate the equivariant Ehrhart series of a lattice polytope and the equivariant Hilbert series of a unimodular triangulation under suitable hypotheses, in particular proving B. Section 6 is devoted to the study of the equivariant Ehrhart series of order polytopes under an action of a finite group via poset automorphisms and contains C among other results. Finally, in Section 7 we apply our results to alcoved polytopes, proving in particular D, and carry out some explicit computations in the case of Lipschitz polytopes of posets.
Disclaimer.
An earlier version of this manuscript proved that (3) holds under some additional hypothesis on the action $\rho$, see Remark 5.5. We subsequently learned that the more general version stated above was proved in unpublished notes by Katharina Jochemko, Lukas Katthän and Victor Reiner. We thank them for sharing their unpublished work and allowing us to include their proof.
Moreover, as the present paper was in preparation, a preprint by Alan Stapledon [Sta23] appeared on arXiv. After some friendly exchanges, the updated versions of both our paper and Stapledon’s make use of each other’s results. In particular, Stapledon obtained an even more general form of (3), see [Sta23, Proposition 4.40, Remark 4.41] and Remark 5.4 below.
Acknowledgements. We thank Alan Stapledon for friendly discussions on his latest work [Sta23] and catching some mistakes in an earlier draft, Victor Reiner for sharing his unpublished work with Katharina Jochemko and Lukas Katthän, the anonymous referee and Matt Beck for their valuable comments. We also thank Mariel Supina for introducing us to the topic with her talk at the workshop Combinatorial and Algebraic Aspects on Lattice Polytopes at Kwansei Gakuin University, and we are grateful to the organizers of this workshop for their support. AD is a member of INdAM–GNSAGA and has been partially supported by the PRIN 2020355B8Y grant “Squarefree Gröbner degenerations, special varieties and related topics” and by the INdAM–GNSAGA project “New theoretical perspectives via Gröbner bases”, CUP E53C22001930001.
@@ 2.3. Translative and proper actions
Translative actions on posets arose from the study of matroid structures related to toric arrangements [DR18], and were further investigated in [DD21] (in the context of invariant rings of Stanley–Reisner rings) and in [BD24] (in connection to supersolvability of geometric posets). If $X$ is the poset of faces of a simplicial fan, proper actions are Stembridge’s proper actions [Ste94]. If $X$ is any poset regarded as a small category without loops, any group action in the sense of Bridson and Haefliger [BH99, Definition 1.11] is proper.
@@ 3.1. Representations of finite groups
Let $G$ be a finite group and $\mathbb{K}$ be a field. A representation of $G$ over $\mathbb{K}$ is any homomorphism $\rho:G\to\operatorname{GL}_{d}(\mathbb{K})$. The representation ring $R(G,\mathbb{K})$ (or “Grothendieck ring of representations”) of $G$ over $\mathbb{K}$ is the ring generated by all isomorphism classes of representations of $G$ over $\mathbb{K}$ with addition and multiplication induced by taking direct sum, resp. tensor product of representations, see [AR23, Definition 2.3].
With every representation $\rho$ is associated a character $\chi_{\rho}:G\mapsto\mathbb{K}$ defined by $\chi_{\rho}(g)=\operatorname{trace}(\rho(g))$. Every character is constant on conjugation classes of $G$ and thus is an element of the ring $\operatorname{C}_{\mathbb{K}}(G)$ of class functions. The ring of virtual characters $\operatorname{R}_{G}(\mathbb{K})$ is the subring of $\operatorname{C}_{\mathbb{K}}(G)$ generated by all integer multiples of characters. The function $\rho\mapsto\chi_{\rho}$ maps $R(G,\mathbb{K})$ to $\operatorname{R}_{G}(\mathbb{K})$. When $\mathbb{K}$ has characteristic zero this mapping is injective and defines an isomorphism onto its image (see, e.g., [CR06, Corollary 30.14]).
@@ 3.2. The equivariant Hilbert series of a Cohen–Macaulay standard graded algebra
where $R_{i}]$ is the isomorphism class of the $G$-representation on $R_{i}$ induced by $\rho$, as an equivariant analogue of the Hilbert series of $R$. Since for the scope of this paper we are only interested in complex representations of finite groups, unless otherwise stated we will assume that $G$ is finite, $\mathbb{K}=\mathbb{C}$, and reserve the name equivariant Hilbert series of $R$ for the following formal power series in $\operatorname{R}_{G}[[t]]$:
@@
where $\chi_{R_{i}}$ is the character of the complex $G$-representation of $R_{i}$ induced by $\rho$.
In the non-equivariant case (over an arbitrary field $\mathbb{K})$, it is customary to express the Hilbert series rationally as $\frac{h(t)}{(1-t)^{d}}$, where $h(t)$ is a polynomial with integer coefficients and $d$ is the Krull dimension of $R$. Such a rational expression is unique and $h(t)$ is known as the $h$-polynomial of $R$. However, when dealing with equivariant Hilbert series, things become more complicated. In the special case when $R$ is the Stanley–Reisner ring of a simplicial complex endowed with a proper action (in the sense of Definition 2.2), we will see in Section 4.1 how to use a suitable multigraded structure of $R$ to obtain a rational expression of the equivariant Hilbert series with a power of (1t)(1-t) as its denominator. Even in that case, though, other rational expressions are possible, and it is not a priori obvious which choice of a denominator is the most informative one.
Recall from commutative algebra that a homogeneous system of parameters for the $\mathbb{N}$-graded $d$-dimensional $\mathbb{K}$-algebra $R$ is a collection $\theta_{1},\ldots,\theta_{d}$ of homogeneous elements of $\mathfrak{m}=(x\in R\mid\deg(x)>0)$ such that $\sqrt{(\theta_{1},\ldots,\theta_{d})}=\mathfrak{m}$. If $\theta_{1},\ldots,\theta_{d}$ all have the same degree $r$, we say that the homogeneous system of parameters has degree $r$; in the special case when $r=1$, we say that the linear forms $\theta_{1},\ldots,\theta_{d}$ form a linear system of parameters for $R$. If $z_{1},\ldots,z_{r}\in R$ are such that $R/(z_{1},\ldots,z_{r})\neq 0$ and, for every $i$, the element $z_{i}$ is a nonzerodivisor on the quotient ring $R/(z_{1},\ldots,z_{i-1})$, we say that the ordered sequence $z_{1},\ldots,z_{r}$ is a regular sequence for $R$. When the ring $R$ is Cohen–Macaulay, then every homogeneous system of parameters is a regular sequence, and this is in fact an alternative characterization of Cohen–Macaulayness for $R$: see, e.g., [Sta96, Theorem I.5.9].
@@
Now take $\ell^{\prime}=x+y$ and $J^{\prime}=(\ell^{\prime})$. Now the denominator is easier to compute; indeed, since $\ell^{\prime}$ is fixed by the $\mathbb{Z}_{2}$-action, we just get that $\det(I-J^{\prime}_{1}t)=1-t$. Since the nontrivial element of $\mathbb{Z}_{2}$ sends $\overline{x}\in R/J^{\prime}$ to $\overline{x}$, the equivariant Hilbert series of $R/J^{\prime}$ is $1+\chi_{\operatorname{sgn}}t$, and overall we have that
@@
This example, despite its simplicity, is quite instructive: indeed, $\{\ell^{\prime}\}=\{x+y\}$ is a linear system of parameters for $R$ that is contained in the invariant subring $R^{\mathbb{Z}_{2}}$. We will see in Theorem 4.6 a characterization of when such linear systems of parameters can be constructed for Stanley–Reisner rings. On the other hand, choosing the linear system of parameters $\{\ell\}=\{x-y\}$ yields a palindromic numerator in the rational expression of the equivariant Hilbert series of $R$, mirroring what happens for the non-equivariant Hilbert series of the Gorenstein ring $R$.